OpenStudy (anonymous):

a motor boat move 18 mph in still water. it travels 35 miles upstream and 35 miles downstream in total time of 4.0 hours. what is the speed of the current?

OpenStudy (anonymous):

this is a bit long but it is step by step time = distance/speed over the ground time upstream = distance/ (boatspeed - current) time downstream = distance/(boatspeed + current) time overall = distance/(boatspeed - current) + distance/(boatspeed + current) Now we need to trun this around to find the current Start by substitute in the numbers we know and replace current with x 4 = 35/(18-x) + 35/(18+x) Now get rid of the fractions multiply both sides by (18-x) 4 (18 -x) = 35 + 35(18-x)/(18+x) multiply both sides by (18+x) 4 (18 -x)(18+x) = 35(18+x) + 35(18-x) multiply out 4 (324 - 18 x + 18x -x^2) = 630 + 35x + 630 - 35x The x's cancel out leaving 4 (324 - x^2) = 1260 multiply out again 1296 - 4 x^2 = 1260 -1260 from both sides 1296 - 1260 -4x^2 = 0 36 - 4x^2 = 0 add 4 x^2 to both sides 36 = 4x^2 9 = x^2 so x=3